# Scalable MHz-Rate Entanglement Distribution in Low-Latency Quantum Networks Interconnecting Heterogeneous Quantum Processors

### v1

@@ I Introduction
Figure 1: An example of quantum data center scheme based on Clos network topology. Three different entanglement distribution scenarios are illustrated: intra-rack, neighboring-rack, and cross-data-center.
@@
Figure 2: a. The conceptual illustration shows the current quantum network solution for heterogeneous entanglement swapping. b. The conceptual illustration of our scalable, programmable solution for distributing entanglement between heterogeneous quantum information processors, which includes quantum computers, memories and sensors. WSS stands for wavelength-selective switch.
@@
Although the successful implementation of QFC has been demonstrated in many experiments, the single-frequency-to-single-frequency conversion protocol has so far failed to allow multiplexing to achieve a higher Bell pair rate. This fact results in the third challenge: an entanglement distribution rate in a single channel stays at the 100 kHz level. Due to the narrow lineshapes of most transitions in matter qubits, the entangled photonic qubits usually have a broad temporal wavepacket [58, 101, 32]. In order to avoid inter-symbol interference between adjacent photonic qubits, the photonic qubit rate for a single channel is thus limited at 100 kHz level [58, 101, 32], resulting in a limited Bell swapping rate. This Bell pair distribution rate is not sufficient to execute error-corrected distributed quantum circuits, as the two-qubit gate execution time can be sub-$\mu$s level [108, 77, 48], and error correction usually consumes multiple Bell pairs for entanglement distillation. Hence, a quantum interface that enables a higher entanglement distribution rate, either through spatial or wavelength multiplexing, is strongly desired, and some early studies indicate possible multiplexing with a limited bandwidth [88].
Unfortunately, the lack of low-loss, low-crosstalk, high-speed, high-dimensional, and large-bandwidth optical switches limits the number of QPUs can be interfaced per network node, further making spatial multiplexing impractical. This challenge comes mainly from the trade-off in optical switches design. Traditional mechanical switches can offer low insertion loss, low crosstalk, large bandwidth, and high-dimensional switching at a cost of low speed and large form factor [73, 46, 66, 121]. For comparison, photonic switches can offer ps-to-nanosecond level switching with a significantly reduced footprint, enabling a fast execution of distributed quantum circuits. However, loss and crosstalk increase significantly as the number of ports of the switch increases [100, 13, 17]. Moreover, photonic switches usually only support a narrow bandwidth of a few nanometers, which is not suitable for the implementation of dense wavelength division multiplexing (DWDM) to boost the entanglement distribution rate. Therefore, a low-loss and low-crosstalk switch with fast reconfiguration speed is still missing for distributed quantum computing (DQC) applications.
@@ II RQI architecture
Figure 3: a. The schematic of the RQI system based on ${{}^{87}\text{Rb}}$ atoms. Atoms are loaded into microcavities via shuttling. The frequency and temporal mode of entangled photons after each microcavity are programmed to the desired frequency and temporal mode to match the photon used in the Bell swapping procedure. All entangled photons are MUXed via a multiplexer to fully leverage the rate advantage of the standard DWDM protocol. The WSS is used to improve the network reconfigurability which is discussed in more details in Section. V. b. The top right figure illustrates the generation of atom-photon entanglement pair using TDM approach. In each microcavity, each atom is prepared and excited sequentially, resulting in the optimal single channel atom-photon entanglement rate.
@@
The programming of the nonlinear process, either a $\chi^{(2)}$ or $\chi^{(3)}$ process, is usually achieved by tuning the refractive index of the material and the spectrum of the pump laser. The tuning mechanism involves thermal, electrical, optical, and optomechanical tuning with a speed ranging from ns to ms levels, which has been studied and applied to various applications in photonics ranging from visible to microwave bands [119, 112, 71, 118, 82, 24, 110, 50, 113, 85, 65, 10, 42, 1, 92, 109, 25, 116, 114]. The on-chip tunable pump laser(s) can be achieved by heterogenous or hybrid integration of the gain medium with tuning realized by photonic integrated circuits [68, 62]. Therefore, reconfiguring the QFC process on integrated photonic circuits is technically available.
@@ III Simulation results of entanglement distribution rates and fidelity
Figure 4: a. Entanglement distribution rate for different DQC scenarios without QFC, with single channel QFC and with DQI DWDM. b. Fidelity of distributed photonic qubits as a function of the number of nodes in the network. Entangled photons are assumed to be emitted at optimal rate from each QPU. After coupled into a SMF, photons travels through the RQI, switches, DWDM multiplexer/demultiplexer and finally the Bell state measurement module. The fiber length, number of switches and DWDM multiplexer/demupltiplexer, as well as the switch latency are different for each scenario with or without RQI. Details of the simulation pipeline can be found in Supplement.
@@ IV RQI design
As mentioned above, the noise of our approach comes mainly from RQI noise, DWDM multiplexer/demultiplexer crosstalk, and switch crosstalk. Except for the RQI noise, the rest of the noise can be modeled from the results in literature [91, 73, 46, 66, 100, 13, 17, 35, 74]. Therefore, our noise and fidelity analysis mainly focuses on the simulation of RQI noise. In this section, we mainly focus on the noise analysis and improvement of the frequency conversion part of the RQI as the relevant research on quantum temporal mode conversion is limited. The diagram of the wavelengths and noise spectrum in our discussion is shown in Fig. 5(a), including the most widely used single-stage $\chi^{(2)}$ difference frequency generation (DFG) process, the low-noise $\chi^{(3)}$ converter based on a four-wave mixing Bragg grating (FWM-BG) process, and the theoretical noise-free $\chi^{(3)}$ converter based on third-order difference frequency generation (TDFG) process.
@@ V Improved network reconfigurability
Figure 6: The schematic of a wavelength selective switch (WSS). Each photon passes through two DWDM multiplexer/demultiplexers and one $N\times M$ switch. The $N\times M$ switch provides all-to-all connectivity, and the $M$ output ports are connected to different DWDM multiplexers. In this example, the $N\times M$ mechanical switch connects ports shown above during the major configuration. In the minor configuration, when the entanglement needs to be distributed from output port 1 to output port 3, the reconfiguration can be rapidly finished by replacing the red DWDM channel with blue DWDM channel.
@@ VI Discussion and conclusion
Trapped ions constitute a major platform for quantum computing, featuring the highest gate fidelity and all-to-all connectivity. However, scaling up the size of a trapped-ion quantum computer is very challenging, and connecting several QPUs within a quantum network is considered as a promising solution. Similarly to neutral atoms, trapped-ion qubits cannot be optically accessed using photons in the telecommunication bands, and thus RQI methods can usefully be employed [96, 15, 53, 31, 89, 52]. The shuttling process in our example can be simplified by assigning each ion chain to an RQI module.
@@
Quantum dots are considered to be a promising candidate for deterministic single-photon emission for many quantum information applications. However, fabricating identical quantum dots remains very challenging, and the wavelength is usually randomly distributed around the desired wavelength because of spectral diffusion. To interfere photons from multiple quantum dots, QFC needs to be implemented. Therefore, with RQI, one can easily multiplex and Bell swap photons emitted by spectrally distinguishable dots without the concern of fabrication defects.
In summary, we propose a quantum networking architecture based on RQI to improve the entanglement distribution rate and reduce the reconfiguration latency. Using the industrial standard DWDM protocol, the entanglement distribution rate can be improved to be greater than 4.5 MHz, with a potential of over 40 MHz, at a reconfiguration time down to the nanosecond level. We analyze the infidelity induced by noisy $\chi^{(2)}$ converters and propose a theoretically noise-free $\chi^{(3)}$ converter design. To the best of our knowledge, this is the first architecture to satisfy the entanglement distribution rate requirement for distributed quantum computing and error correction. Our approach is based on the standard DWDM infrastructure, which indicates a promising low-cost solution to enable distributed quantum computing.
@@ Appendix A Intra-cavity TDM
Recently, a new research has mentioned that it is possible to have a higher attempt rate and a lower initialization time for atoms [93]. After adopting the new parameters, a higher upper bound of entanglement rate between two atoms is found to be 2.3 MHz.
@@ Appendix B Rate and fidelity simulation pipeline
Now we discuss how we calculate $P_{suc}$ in our rate simulation. We assume a photon collection efficiency $\eta_{col}=0.96$ and a SMF coupling efficiency $\eta_{coup}$ of 0.9, which is consistent with previous work [58, 101, 115]. The typical fiber loss for DWDM channels is $l_{tele}=0.17$dB/km (Corning SMF-28), and for 780 nm, the loss $l_{NIR}$ is usually around 4dB/km (Corning HI780). Another major loss is the insertion loss from switches and DWDM multiplexers/demultiplexers. A designated quantum matrix switch can have a typical insertion loss of 0.35 dB and the worst insertion loss of 0.7 dB for DWDM channels [35], while such a switch in NIR can have a loss of more than 2 dB. A much higher-dimension switch is also available with a similar performance in terms of loss (worst loss less than 1 dB for the 192$\times$192 switch) [20]. Note that, the crosstalk for these mechanical-based optical switches is usually less than 70 dB, which is significantly lower than the photonics counterparts. The switch loss 1.33 dB used by previous work only supports telecommunication bands with mechanical switches [58, 78, 73, 98]. For photonic switches, the state-of-art loss is found to be around 4 dB for a 2 by 8 switch [2]. Considering that the technology for photonic switches is growing rapidly, we use a 2 dB loss for photonic switches for even high-dimensional switches in our simulation.
@@ Appendix C Raman noise analysis in PPLN waveguides
Fig. 7(a) and 7(b) show the temperature-dependent peak frequency and the line width of the three dominant optical phonon modes in an x-cut 5% MgO:LN crystal. The peak frequency undergoes a slight redshift, and the line width increases as the temperature of the system rises. Such temperature-dependent variation gives rise to a noticeable change in the Raman susceptibility curve whose imaginary part is shown in Fig. 7(c).
Considering the temperature dependence of the Raman susceptibility, we can now determine the noise spectrum density $\frac{dN_{s}}{d\lambda}$ as a function of the Raman shift $\Omega$ for a pump power $P_{0}=200$ mW and a waveguide length of 3 cm. The anti-Stokes scattering is much weaker than the Stokes scattering because of the difference in the Bose-Einstein population factor, so a long-wavelength pump scheme (the wavelength of the pump is the largest among all three waves being mixed) is usually adopted in common QFC devices. Therefore, only the anti-Stokes scattering is considered as the main source of Raman noise in the main text. The curves in Fig. 8 in the supplementary material show the noise spectral density of the Stokes photons under different temperatures for a target signal wavelength at 1.52 $\mu$m. As the pump wavelength moves further away from the target signal wavelength, the Stokes scattering is dramatically suppressed. However, compared to the anti-Stokes scattering shown in Fig. 5(b) of the main text, the Stokes scattering noise is, on average, one to two orders of magnitude higher than the anti-Stokes noise, and the anti-Stokes component becomes more sensitive to the temperature change at longer pump wavelengths. Hence, the difference frequency generation with a long-wavelength pump is well justified in our QFC application.

### v2

@@ I Introduction
Figure 1: a. The conceptual illustration shows the current quantum network solution for heterogeneous entanglement swapping. The four existing challenges on different components of the quantum network are shown correspondingly. b. The conceptual illustration of our scalable, programmable solution for distributing entanglement between heterogeneous quantum information processors, which includes quantum computers, memories and sensors. WSS stands for wavelength-selective switch. The detailed configuration of our system is shown in Fig. 2
@@
Although the successful implementation of QFC has been demonstrated in many experiments, the single-frequency-to-single-frequency conversion protocol has so far failed to allow multiplexing to achieve a higher Bell pair rate. This fact results in the third challenge: an entanglement distribution rate in a single channel stays at the 100 kHz level. Due to the narrow lineshapes of most transitions in matter qubits, the entangled photonic qubits usually have a broad temporal wavepacket [7, 8, 9]. In order to avoid inter-symbol interference between adjacent photonic qubits, the photonic qubit rate for a single channel is thus limited at 100 kHz level [7, 8, 9], resulting in a limited Bell swapping rate. This Bell pair distribution rate is not sufficient to execute error-corrected distributed quantum circuits. Executing high-fidelity remote gates in a timely manner requires the MHz rate entanglement distribution, as the two-qubit gate execution time can be sub-$\mu$s level [50, 51, 52], and high-fidelity Bell pairs from purification usually consume multiple noisy Bell pairs. Hence, a quantum interface that enables a higher entanglement distribution rate, either through spatial or wavelength multiplexing, is strongly desired, and some early studies indicate possible multiplexing with a limited bandwidth [53].
Unfortunately, the lack of low-loss, low-crosstalk, high-speed, high-dimensional, and large-bandwidth optical switches limits the number of QPUs can be interfaced per network node, further making spatial multiplexing impractical. This challenge comes mainly from the trade-off in optical switches design. Traditional mechanical switches can offer low insertion loss, low crosstalk, large bandwidth, and high dimensional switching at a cost of low speed and large form factor [10, 11, 12, 54]. For comparison, photonic switches can offer ps-to-nanosecond level switching with a significantly reduced footprint, enabling a fast execution of distributed quantum circuits. However, loss and crosstalk increase significantly as the dimension (number of ports on the switch) increases [13, 14, 55]. Moreover, photonic switches usually only support a narrow bandwidth of a few nanometers, which is not suitable for the implementation of dense wavelength division multiplexing (DWDM) to boost the entanglement distribution rate. Besides, a slow reconfiguration time of switch can significantly reduce the entanglement distribution rates, introducing non-negligible latency in remote quantum circuit execution [90]. Therefore, a low-loss and low-crosstalk switch with fast reconfiguration speed is still missing for distributed quantum computing (DQC) applications.
@@ II RQI architecture
Figure 2: a. The schematic of the RQI system based on ${{}^{87}\text{Rb}}$ atoms. Atoms are loaded into microcavities via shuttling. The frequency and temporal mode of entangled photons after each microcavity are programmed to the desired frequency and temporal mode to match the photon used in the Bell swapping procedure. All entangled photons are MUXed via a multiplexer to fully leverage the rate advantage of the standard DWDM protocol. The WSS is used to improve the network reconfigurability which is discussed in more details in Section. V. b. The top right figure illustrates the generation of atom-photon entanglement pair using TDM approach. In each microcavity, each atom is prepared and excited sequentially, resulting in the optimal single channel atom-photon entanglement rate. The design and realization of RQI can be found in Section. IV and Fig. 4
@@
The programming of the nonlinear process, either a $\chi^{(2)}$ or $\chi^{(3)}$ process, is usually achieved by tuning the refractive index of the material and the spectrum of the pump laser. The tuning mechanism involves thermal, electrical, optical, and optomechanical tuning with a speed ranging from ns to ms levels, which has been studied and applied to various applications in photonics ranging from visible to microwave bands [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. The on-chip tunable pump laser(s) can be achieved by heterogenous or hybrid integration of the gain medium with tuning realized by photonic integrated circuits [87, 88]. Therefore, reconfiguring the QFC process on integrated photonic circuits is technically available.
@@ III Simulation results of entanglement distribution rates and fidelity
Figure 3: a. An example of quantum data center based on Clos network topology. Three different distributed quantum computing scenarios are illustrated: intra-rack, neighboring-rack, and cross-data-center. b. Entanglement distribution rate for single channel without QFC, single channel QFC and DQI DWDM. c. Fidelity of distributed photonic qubits as a function of the number of nodes in the network. Entangled photons are assumed to be emitted at optimal rate from each QPU. Details of the simulation pipeline can be found in Appendix B.
@@ IV RQI design
As mentioned above, the noise of our approach comes mainly from RQI noise, DWDM multiplexer/demultiplexer crosstalk, and switch crosstalk. Except for the RQI noise, the rest of the noise can be modeled from the results in literature [92, 10, 11, 12, 13, 14, 55, 93, 94]. Therefore, our noise and fidelity analysis mainly focuses on the simulation of RQI noise. In this section, we mainly focus on the noise analysis and improvement of the frequency conversion part of the RQI as the relevant research on quantum temporal mode conversion is limited. The diagram of the wavelengths and noise spectrum in our discussion is shown in Fig. 4(a), including the most widely used single stage $\chi^{(2)}$ difference frequency generation (DFG) process, the low-noise $\chi^{(3)}$ converter based on a four-wave mixing Bragg grating (FWM-BG) process, and the theoretical noise-free $\chi^{(3)}$ converter based on third-order difference frequency generation (TDFG) process.
@@ V Improved network reconfigurability
Figure 5: The illustration of two types of reconfiguration using RQI array and a wavelength selective switch (WSS). Each photon passes through two DWDM multiplexer/demultiplexers and one $N\times M$ switch. The $N\times M$ switch provides all-to-all connectivity, and the $M$ output ports are connected to different DWDM multiplexers. In this example, the $N\times M$ mechanical switch connects ports via solid red lines, and the user needs to reroute the photon from output port 1 to output port 3. If the minor configuration is used, the reconfiguration can be rapidly finished by tuning the photon from the red DWDM channel to blue DWDM channel at GHz level. If the major reconfiguration is used, the $N\times M$ mechanical switch reconfigure the connection from solid red line to dashed red line at kHz rate.
@@ VI Discussion and conclusion
Trapped ions constitute a major platform for quantum computing, featuring the highest gate fidelity and all-to-all connectivity. However, scaling up the size of a trapped-ion quantum computer is very challenging, and connecting several QPUs within a quantum network is considered a promising solution. Similarly to neutral atoms, trapped-ion qubits cannot be optically accessed using photons in the telecommunication bands, and thus RQI methods can be used in a useful way [17, 18, 19, 20, 21, 22]. The shuttling process in our example can be simplified by assigning each ion chain to an RQI module.
@@
Quantum dots are considered to be a promising candidate for deterministic single-photon emission for many quantum information applications. However, fabricating identical quantum dots remains a challenge, and the wavelength is usually randomly distributed around the desired wavelength because of spectral diffusion. To interfere photons from multiple quantum dots, QFC needs to be implemented. Therefore, with RQI, one can easily multiplex and Bell swap photons emitted by spectrally distinguishable dots without the concern of fabrication defects.
Table 1: A comparison of distributed quantum computing metrics between our RQI architecture and current solutions. Rate represents the entanglement distribution rate.
Single channel without QFC Single channel with QFC RQI DWDM
Intra-rack Rate (kHz) [^1] Assume use GHz photonic switch. 25.16 24.66 4508
Inter-rack Rate (kHz) 9.66 13.34 3844
Intra-rack Rate (kHz) 0.040 5.96 2696
Fidelity [^2] Assume perfect atom-photon entanglement. (3 nodes) 0.987 0.938 0.946
Fidelity [^2] Assume perfect atom-photon entanglement. (9 nodes) 0.924 0.878 0.926
[…]
Connectivity All-to-all All-to-all All-to-all [^3] All-to-all connectivity in major reconfiguration.
In summary, we propose a quantum networking architecture based on RQI to improve the entanglement distribution rate and reduce the reconfiguration latency. As summarized in Table 1, compared to current quantum network architectures, our architecture can improve the entanglement distribution rate to be greater than 4.5 MHz, with a potential of more than 40 MHz, using the industrial standard DWDM protocol. By dropping inessential connectivity within the circuit execution, the reconfiguration time can be reduced down to the nanosecond level without sacrificing the entanglement distribution rate and fidelity. In terms of the RQI design, we analyze the infidelity induced by noisy $\chi^{(2)}$ converters and propose a theoretically noise-free $\chi^{(3)}$ converter design. To the best of our knowledge, this is the first architecture to satisfy the entanglement distribution rate requirement for distributed quantum computing and error correction. Our approach is based on the standard DWDM infrastructure, which indicates a promising low-cost solution to enable distributed quantum computing.
@@ Appendix A Intra-cavity TDM
Recently, a new research has mentioned that it is possible to have a higher attempt rate and a lower initialization time for atoms [6]. After adopting the new parameters, a higher upper bound of the entanglement rate between two atoms is found to be 2.3 MHz.
@@ Appendix B Rate and fidelity simulation pipeline
Now we discuss how we calculate $P_{suc}$ in our rate simulation. We assume that the ${{}^{87}\text{Rb}}$ atom has a photon emission in cavity probability of 0.5. The collection efficiency for this photon $\eta_{col}=0.96$ and a SMF coupling efficiency $\eta_{coup}$ of 0.9, which is consistent with previous work [7, 8, 56]. The typical fiber loss for DWDM channels is $l_{tele}=0.17$dB/km (Corning SMF-28), and for 780 nm, the loss $l_{NIR}$ is usually around 4dB/km (Corning HI780). Another major loss is the insertion loss from switches and DWDM multiplexers/demultiplexers. A designated quantum matrix switch can have a typical insertion loss of 0.35 dB and the worst insertion loss of 0.7 dB for DWDM channels [93], while such a switch in NIR can have a loss of more than 2 dB. A much higher-dimension switch is also available with a similar performance in terms of loss (worst loss less than 1 dB for the 192$\times$192 switch) [109]. Note that, the crosstalk for these mechanical-based optical switches is usually less than 70 dB, which is significantly lower than the photonics counterparts. The switch loss 1.33 dB used by previous work only supports telecommunication bands with mechanical switches [7, 110, 10, 91]. For photonic switches, the state-of-art loss is found to be around 4 dB for a 2 by 8 switch [111]. Considering that the technology for photonic switches is growing rapidly, we use a 2 dB loss for photonic switches for even high-dimensional switches in our simulation.
@@ Appendix C Raman noise analysis in PPLN waveguides
Fig. 6(a) and 6(b) show the temperature-dependent peak frequency and the line width of the three dominant optical phonon modes in an x-cut 5% MgO:LN crystal. The peak frequency undergoes a slight redshift, and the line width increases as the temperature of the system rises. Such temperature-dependent variation gives rise to a noticeable change in the Raman susceptibility curve whose imaginary part is shown in Fig. 6(c).
Considering the temperature dependence of the Raman susceptibility, we can now determine the noise spectrum density $\frac{dN_{s}}{d\lambda}$ as a function of the Raman shift $\Omega$ for a pump power $P_{0}=200$ mW and a waveguide length of 3 cm. The anti-Stokes scattering is much weaker than the Stokes scattering because of the difference in the Bose-Einstein population factor, so a long-wavelength pump scheme (the wavelength of the pump is the largest among all three waves being mixed) is usually adopted in common QFC devices. Therefore, only anti-Stokes scattering is considered as the main source of Raman noise in the main text. The curves in Fig. 7 in the Appendix show the noise spectral density of the Stokes photons under different temperatures for a target signal wavelength at 1.52 $\mu$m. As the pump wavelength moves further away from the target signal wavelength, the Stokes scattering is dramatically suppressed. However, compared to the anti-Stokes scattering shown in Fig. 4(b) of the main text, the Stokes scattering noise is, on average, one to two orders of magnitude higher than the anti-Stokes noise, and the anti-Stokes component becomes more sensitive to the temperature change at longer pump wavelengths. Hence, the difference frequency generation with a long-wavelength pump is well justified in our QFC application.